The enrollment at a school in 1998 was 55. In 2004, the enrollment was 145. Fine the average rate of change in enrollment from 1998 to 2004.: The enrollment at a school in 1998 was 55. In 2004, the enrollment was 145. Fine the average rate of change in enrollment from 1998 to 2004.
Answer by checkley77(3654)

(145-55)/6=90/6=15 is the average enrollment increase per year.